# FedDRL: A Trustworthy Federated Learning Model Fusion Method Based on Staged Reinforcement Learning

### 21 Jan 2024 v3

@@ 5.2.2 Low-quality model fusion experiments
To validate the performance of our proposed algorithm FedDRL, we performed validation on three datasets, Fashion-MNIST, CIFAR-10, and CIFAR-100. Since all three are open source, these data are high-quality data. Therefore each client used these sample data to train their respective models, and these models did not exhibit significant accuracy differences. We set up the following experiments to verify the impact of low-quality models involved in global model fusion in a real-world scenario.
(1) To verify the impact of different numbers of customer nodes on the algorithm, we conducted experiments on three datasets with 5, 10, and 15 clients, respectively. We also set the parameter alpha=1 of Dirichlet to complete the dataset partitioning corresponding to each client.
(2) We introduced a method for low-quality models to participate in the fusion. At the beginning of each round of federated learning, these clients initialize the local model using the global model. After several rounds of communication, when the accuracy of the local model reaches a certain accuracy, we simulate the participation of low-quality models in global model fusion by controlling the accuracy of these client models (e.g., model accuracy is kept between 50% and 60%). These clients upload low-quality models to the server side. Our experimental configuration of low-quality simulation for a different number of clients in table 3.
@@
Client=5 Fashion-MINST Client1 9061 53%
@@
Client=10 Fashion-MINST Client1,Client5 5071, 7245 51%, 52%
@@
Client=15 Fashion-MINST Client1,Client5,Client10 4405, 3752, 1809 52%, 51%, 53%
@@
We compared the FedDRL algorithm with both FedAvg and FedProx algorithms. We set the number of communication rounds to 100. Each client performs one epoch of local model training. The experimental results are shown in table 4.
Table 4: Global model accuracy of Low-quality model experiments
@@
Clients=5 Clients=10 Clients=15 Clients=5 Clients=10 Clients=15 Clients=5 Clients=10 Clients=15
@@
To demonstrate the FedAvg, FedProx, and FedDRL algorithms for global model fusion at each communication round. We have taken the CIFAR10 dataset as an example and conducted experimental comparisons with 5, 10, and 15 clients, respectively. The experimental results of each algorithm are shown in Figure 5.
The comparison experiments of CIFAR-10 show that the performance of FedAvg and FedProx algorithms keeps decreasing as the number of low-quality models involved in fusion increases. Also, these two algorithms fluctuate up and down with the accuracy of the low-quality models due to the number of samples used by these two algorithms to calculate the global model. Therefore, when the weight value of the model is fixed, the change in the accuracy of the model affects the change in the accuracy of the global model. Meanwhile, FedDRL has a faster convergence speed than both FedAvg and FedProx algorithms because the FedDRL algorithm can adjust the corresponding weights adaptively according to the quality of each client model, which mitigates the impact of low-quality models on the accuracy of the global model and thus accelerates the convergence speed of the global model to a certain extent.
Refer to caption (a) 5 clients FedAvg
Refer to caption (b) 5 clients type2 FedProx
Refer to caption (c) 5 clients FedDRL
Refer to caption (d) 10 clients FedAvg
Refer to caption (e) 10 clients FedProx
Refer to caption (f) 10 clients FedDRL
Refer to caption (g) 15 clients FedAvg
Refer to caption (h) 15 clients FedProx
Refer to caption (i) 15 clients FedDRL
Figure 7: Global model accuracy in Low-quality model.
@@ 5.2.3 Hybrid experiment
In this section, we fuse two attack types (type 1 and type 3) and simulate a low-quality model to participate in the fusion. We verify The performance of the FedDRL algorithm in this hybrid Scenario. We also compare with FedAvg and FedProx methods.
We used the cifar10 dataset and conducted experiments with 10 and 15 client cases. Client 1 only uploads the initialization model for each training, Client 6 simulates low-quality models to participate in fusion, client 10 or Client 11 initiates attack behavior in odd-numbered communication rounds, and even-numbered rounds participate in training normally. The remaining nodes normally participate in each round of Federated learning tasks. The detailed configuration of the experiments is shown in table 5. After 100 communication rounds, the global model accuracy of each algorithm is shown in table 6.
Table 5: Hybrid Experimental Settings for Different Numbers of Clients
@@
Client=10 Client1 Attack Type 1 4222 $A\leq 10\%$
@@
Client=15 Client1 Attack Type 1 3670 $A\leq 10\%$
@@
Table 6: Global model Accuracy of Hybrid Experimental in different clients’ number
@@
Refer to caption (a) 10 clients FedAvg
Refer to caption (b) 10 clients FedProx
Refer to caption (c) 15 clients FedDRL
Refer to caption (d) 15 clients FedAvg
Refer to caption (e) 15 clients FedProx
Refer to caption (f) 15 clients FedDRL
The hybrid scenario experiments show that the FedAvg and FedProx algorithms can no longer perform global model fusion properly due to the attack behavior of the client. Meanwhile, malicious models’ involvement in the global model fusion makes the accuracy of the global model based on the traditional algorithm very low. In the FedDRL algorithm, we adopt the two-stage method. In the first stage, we eliminate malicious models from participating in a fusion. We use an adaptive weight fusion strategy in the second stage, reducing low-quality models’ influence. Thus our algorithms work correctly even in this hybrid scenario. The global model and the model accuracy of each client for each communication round of these three algorithms are shown in Figure 8.
5.2.4 Discussion
We analyze the experimental results for different model fusion scenarios as well as for different numbers of client models. In the client-side attack experiments, under attack type 1, our algorithm works better than the FedAvg method on all three datasets. Meanwhile, as the number of malicious nodes increases, the global model accuracy of both FedAvg and FedProx algorithms shows a serious decrease. In contrast, our algorithm still maintains a smooth convergence trend. Under attack type 2, our algorithm is better than the FedAvg algorithm and slightly lower than the FedProx algorithm. Under attack type 3, our algorithm outperforms the edAvg and FedProx algorithms on all three datasets, which no longer work under rotating attacks with multiple malicious clients. In contrast, our algorithm can work reliably. In the low-quality model experiments, our algorithm outperforms the FedAvg and FedProx algorithms on all three datasets. In the hybrid experiments, the FedAvg and FedProx algorithms no longer work properly, and our algorithm can fuse the global models properly.
Since the FedAvg and FedProx methods use the number of samples to calculate the model weights, they cannot adaptively adjust the responses of malicious and low-quality models. They thus cannot be applied to these attack scenarios. Our proposed method is based on staged reinforcement learning for malicious node removal and optimal weight assignment for trusted clients and thus can work reliably in various environments.
@@ 6 CONCLUSION
To solve the problem of assigning client weights due to data heterogeneity in a federated learning environment. We propose FedDRL, an adaptive weight calculation method based on staged reinforcement learning. The method comprises two stages: selecting trusted clients and adaptive weight assignment. In the first stage, our approach can eliminate malicious client models from participating in the global model fusion. It also can dynamically select trusted client models to participate in the next stage of model weight assignment. In the second stage, Our method can adaptively calculate the corresponding weights based on the model quality of each client. Also, we propose a distributed reinforcement learning method to accelerate the agent’s training. Finally, we proposed five types of model fusion scenarios to validate our approach, and experiments show that our proposed algorithm can work reliably in all kinds of model fusion scenarios while maintaining the accuracy of the global model.
Currently, our algorithm requires separate training of the agent according to the client fusion scenarios. In future work, we will extend the proposed algorithm to accommodate model fusion tasks for different task scenarios. We will also continue improving the algorithm’s performance to achieve more client tasks.

### 19 Mar 2024 v4

@@ 5.2.2 Low-quality Model Fusion Experiments
In evaluating our FedDRL framework, we undertook validation using the Fashion-MNIST, CIFAR-10, and CIFAR-100 datasets. Given their open-source nature, these datasets are of high quality, leading to minimal variance in model accuracy among clients utilizing them directly. Thus, to simulate real-world conditions, we incorporated low-quality models into the global fusion process. We established a model accuracy threshold, ensuring that models uploaded by low-quality clients did not exceed this threshold in any communication round.
Experiments were carried out on the three datasets, with client groups of varying sizes—5, 10, and 15—participating in the global model fusion. We applied a Dirichlet distribution with parameter alpha=1 to achieve dataset segmentation among clients. We set some clients to upload low-quality models; after several communication rounds, we controlled these client models’ accuracy in global fusion, ensuring it remained within the 40% to 55% range.
Details of these low-quality model experiment configurations are specified in Table 3. The FedDRL algorithm was compared against the FedAvg and FedProx methods across 100 communication rounds, with each client executing one epoch of local model training. Results are summarized in Table 4.
@@
Clients=5 Fashion-MINST Client1 9061 53%
@@
Clients=10 Fashion-MINST Client1,Client5 5071, 7245 51%, 52%
@@
Clients=15 Fashion-MINST Client1,Client5,Client10 4405, 3752, 1809 52%, 51%, 53%
@@
The FedDRL algorithm was compared against the FedAvg and FedProx methods across 100 communication rounds, with each client executing one epoch of local model training. Results are summarized in Table 4. Details of these low-quality model experiment configurations are specified in Table 3.
Table 4: Accuracy of each algorithm for low-quality modeling experiments
@@
C=5 C=10 C=15 C=5 C=10 C=15 C=5 C=10 C=15
@@
Employing the CIFAR-10 dataset for illustrative purposes, we performed comparative analyses for setups with 10 and 15 clients, respectively; the findings are depicted in Figure 7. The experiments indicate that the accuracy of the FedAvg and FedProx methods deteriorates as the prevalence of low-quality models increases. This decline can be attributed to these algorithms’ reliance on sample count for determining the fusion weight values of the models, where the inclusion of low-quality models adversely impacts the global model’s accuracy. Conversely, FedDRL surpasses both methodologies in terms of global model convergence speed and accuracy. This is because FedDRL adaptively recalibrates the weights assigned to each client’s model based on quality, thereby diminishing the adverse effects of low-quality models on the global model’s accuracy and consequently hastening the global model’s convergence rate.
Refer to caption (a) 10 clients FedAvg
Refer to caption (b) 10 clients FedProx
Refer to caption (c) 10 clients FedDRL
Refer to caption (d) 15 clients FedAvg
Refer to caption (e) 15 clients FedProx
Refer to caption (f) 15 clients FedDRL
Figure 7: The accuracy of a global model for different numbers of client in Low-quality scenario.
@@ 5.2.3 Hybrid experiment
In this section, we establish a hybrid scenario incorporating two types of attacking clients (type 1 and type 3) alongside clients submitting low-quality models. We assess the effectiveness of the FedDRL algorithm within this mixed scenario and benchmark it against the FedAvg and FedProx approaches.
Employing the CIFAR-10 dataset, we set different numbers of clients (10,15) participating in global model fusion, respectively. Client 1 persistently uploads merely the initial model at each round. Client 6 emulates the submission of low-quality models for fusion, and Client 10 or 11 engages in attack behaviour during odd communication rounds but normally participates during even rounds. The remainder of the nodes contribute routinely to each cycle of the federated learning tasks. The experimental setup specifics are delineated in Table 5.
Table 5: Experimental settings for hybrid scenarios
@@
Clients=10 Client1 Attack Type 1 4222 $A\leq 10\%$
@@
Clients=15 Client1 Attack Type 1 3670 $A\leq 10\%$
@@
Refer to caption (a) 10 clients FedAvg
Refer to caption (b) 10 clients FedProx
Refer to caption (c) 10 clients FedDRL
Refer to caption (d) 15 clients FedAvg
Refer to caption (e) 15 clients FedProx
Refer to caption (f) 15 clients FedDRL
After completing 100 communication rounds, we present the global model accuracy for each algorithm in Table 6. The comparative global model accuracies and individual client model accuracies per communication round, as determined by these three algorithms, are depicted in Figure 8. The experimental outcomes from the hybrid scenario reveal that the FedAvg and FedProx algorithms falter in properly conducting global model fusion due to the adversarial behavior of certain clients. Incorporating malicious models under traditional algorithmic frameworks significantly degrades the global model’s accuracy.
Table 6: Accuracy of each algorithm for hybrid scenarios experiments
@@
The experimental outcomes show that FedAvg and FedProx’s global model accuracies suffer from malicious attacks due to their weighted average-based fusion, which doesn’t block harmful participants. Conversely, the FedDRL algorithm, through its two-stage approach, initially filters out malicious models from fusion and subsequently applies an adaptive weight strategy to diminish the impact of substandard models. Consequently, our algorithm maintains operational integrity even within this complex scenario.
5.2.4 Agent Training Efficiency in the FedDRL Framework
In this segment, our primary objective is to assess the training efficiency of agents within the FedDRL framework. To expedite the training process, we have implemented optimizations in two key areas. Initially, we adopted a distributed reinforcement learning methodology, enabling multi-agents to interact concurrently with the external environment. Concurrently, we introduced a memory cache module designed to prevent redundant sampling by multiple agents.
Experimental Scenarios: Our investigation encompasses varied attack scenarios across two distinct datasets: Fashion-MNIST and Cifar-10. In each scenario, we involve a total of 10 and 15 clients in the federated task, including 2 and 3 malicious clients accordingly.
Comparison Experiments: To ascertain the efficacy of the FedDRL framework, we initiated experiments featuring 1, 5, 10, and 20 agents. To guarantee the stability of the reward values acquired by the final agents, we designated the number of iterations for each experimental group to be 10,000, 15,000, 20,000, and 25,000, correspondingly.
Experimental metrics: Our evaluation involves counting the iterations necessary for reinforcement learning to reach stable rewards across different agent counts. We employ a sliding window approach to compute the average reward, depicting the progression of rewards attained by the agents. We define $r_{t}$ the agent’s reward obtains in the $t$-th interaction and the sliding window as $W$. The formula for calculating the average reward is represented as equation 22:
$\bar{R}=\frac{1}{W}{\textstyle\sum_{t=1}^{W}}r_{t}$ (22)
Reward Parameter Setting: Our reward function comprises two components: the global model accuracy reward and the reward for the number of credible nodes. For this experiment, these parameters are set to 100 and 10, respectively.
Table 7: The iterations of obtaining stable rewards for different numbers of agents
Dataset Attack Type The number of agents
N=1 N=5 N=10 N=20
Fashion-MNIST Type-1 25000 20000 16000 8000
Type-2 25000 21000 15000 9000
CIFAR-10 Type-1 25000 20000 12000 10000
Type-2 25000 20000 14000 9000
Refer to caption (a) Fashion-MNIST (Attack Type 1)
Refer to caption (b) Fashion-MNIST (Attack Type 2)
Refer to caption (c) Cifar-10 (Attack Type 1)
Refer to caption (d) Cifar-10 (Attack Type 2)
Figure 9: Global model accuracy in 15 clients attack type 1
Experimental Results: In accordance with our experimental setup, we recorded the reward values for each iteration of the agents, as detailed in Figure 9. We have systematically arranged this information into Table 7 for enhanced clarity regarding the actual iterations across different experiments.
The data reveals a notable trend: The single agent does not get the optimal reward in some attack scenarios, because the single agent is easy to fall into the local optimal solution. Meanwile, an increase in agents correlates with reducing the iterations required to achieve a stable reward. However, this relationship is not strictly proportional because the multi-agent independently train their respective Actor and Critic networks. Each agent necessitates a distinct number of iterations to ensure the stability of its individual networks. Nevertheless, the simultaneous interaction of multiple agents with the environment markedly decreases the sampling time, demonstrating a clear trade-off between computational resources and time. This strategy underlines the significant computational resources required, highlighting a deliberate exchange of increased computational demand for reduced computational time.
@@ 6 CONCLUSION
To realize trustworthy federated learning, We propose a trusted reinforcement learning framework (FedDRL) based on staged reinforcement learning. The framework comprises two phases: selecting trusted clients and adaptive weight assignment. In the first phase, we design a reward strategy to train the agent, which allows the trained agent to exclude malicious client models from participating in model fusion based on the environment, and it also adaptively selects trustworthy clients for model fusion. In the second phase, we design a dynamic model weight calculation method, which can adaptively calculate the corresponding weights based on the model quality of each client. In addition, we propose a distributed reinforcement learning method to accelerate agent training. Finally, we design five model fusion scenarios to validate our approach, and the experiments show that our proposed algorithm can work reliably in various model fusion scenarios while maintaining global model accuracy.
Although a multi-agent distributed reinforcement learning approach can accelerate the agent training process, it sacrifices computational resources for the computational time. In our future work, we will continue to explore more lightweight and trustworthy federated learning methods. We will also investigate more efficient reinforcement learning methods for credible federated learning.